OpenStudy (anonymous):

A steel pipe in the shape of a right cylinder is used for drainage under a road. The length of the pipe is 12 feet and the diameter is 36 inches. The pipe is open at both ends. The outer surface of the pipe is coated with protective material. How many square feet of the protective material will be needed?

OpenStudy (anonymous):

Area of cylinder=2πrh+2πr^2 36 inches=3 feet r=3/2 feet or 1.5 feet =2π(1.5)(12)+2π(1.5)^2 =113.0973355+14.13716694 =127.2345024 (total area of cylinder with 2 ends) because the cylinder has two open ends therefore subtract the 2πr^2 =113.0973355 or 113.097 sq feet Therefore you need 113.097 sq feet of protective material